OpenStudy (boxman61):

write as radical then simplify 4^-1/2

OpenStudy (anonymous):

What's 1/sqrt(4)?

OpenStudy (anonymous):

Box man yours \[4^-1/2 = \sqrt{4} \] which equals 2 Brinethery yours is 1/2

OpenStudy (boxman61):

\[4^{-1/2} = 1/\sqrt{4} = 1/2\]

OpenStudy (boxman61):

thanks

OpenStudy (anonymous):

yw